# Fréchet Power-Scenario Distance: A Metric for Evaluating Generative AI Models across Multiple Time-Scales in Smart Grids

### v1

@@ IV-D Cross-model comparison with FPD
In the previous sections, all evaluations were conducted directly on the artificially generated datasets. This approach enables precise control over generation quality and model performance assessment. However, FPD can also be used to evaluate model generated performing the same task, as demonstrated in the following two scenarios.
@@
    Scenario 1: EV Charging Profile Data Generation
    Accurately evaluating the quality of synthetic data in real-world settings presents a significant challenge, as highlighted by [24], for which FPD provides a solution. [24] proposed a novel generative model, DiffCharge, based on a diffusion model designed to conditionally generate electric vehicle (EV) charging power profiles. Probability density function (PDF) comparisons and discriminative scores derived from a trained classifier were used to evaluate battery-level data generation performance between DiffCharge and baseline models such as Time-GAN. However, evaluations for station-level data generation were not given. We conduct experiments to compare DiffCharge’s performance with a W-GAN model on the station level; see Fig. 7. We find that DiffCharge achieves a lower FPD score. Visually, we also observe that DiffCharge can capture the general trend better, while W-GAN generates less realistic peaky signals. Therefore, our results indeed support the conclusion that the DiffCharge model can more accurately approximate the distribution of EV charging profiles at both the station-level and battery-level.
@@
    Scenario 2: Wind Power Data Generation
    To further demonstrate FPD’s cross-model comparison ability, we replicate the study from [1], where the authors introduced a model called StyleGAN, which integrates forecast and climate data to enhance the generation of renewable energy data. They showed that StyleGAN is superior to WGAN proposed in [23] based on multiple uncertainty quantification and statistical metrics, such as RMSE. In our experiment, the differences between generations from StyleGAN and WGAN were challenging to distinguish through visual inspection. StyleGAN’s results seemed to capture the overall trend more accurately. As shown in Fig. 8, FPD confirms StyleGAN aligns better with real data, supporting the findings in [1].
    Figure 7: Scenario 1 – Comparison of DiffCharge and W-GAN in generating EV charging profiles at the station level.
    Figure 8: Scenario 2 – Comparison of W-GAN and StyleGAN in generating wind power data.
IV-E Comparability across varying time resolutions
A key feature of FPD is its ability to adapt to varying time resolutions within a single metric. This flexibility enables effective comparisons across datasets with different time resolutions. To demonstrate this capability, we collected hourly resolution data for solar and wind power and evaluated their similarity with the 5-minute resolution solar and wind power datasets used in Section IV-A. The results, shown in Fig. 9, indicate that FPD successfully identifies that the solar power datasets are more similar to each other than they are to the wind power data despite their differing time resolutions. This result demonstrates the metric’s effectiveness in facilitating comparability across datasets with different temporal resolutions.
Figure 9: The FPD metric is evaluated between the left:5-min resolution solar power dataset[17], right top:hourly resolution solar power dataset[25] and right bottom:hourly resolution wind power dataset[25]. The results demonstrate that FPD effectively generalizes across different time resolutions.
@@ V Conclusion and Future work
This paper introduces the Fréchet Power-Scenario Distance (FPD) as a novel, task-agnostic metric for evaluating generative model outputs in the power systems sector. FPD addresses the limitations of traditional metrics by employing a hierarchical feature extraction framework that captures multi-resolution temporal information. Then, the Fréchet Distance is applied to assess the quality of generated data in the feature space. This method enables FPD to provide an accurate, robust, and consistent evaluation across diverse power system applications.
Comprehensive case studies demonstrate FPD’s effectiveness in distinguishing data quality across various disturbances, time resolutions, and sub-task scenarios. These results underscore FPD’s potential as a standardized cross-model approach for validating generative models, offering support for both enhancing model quality and promoting wider adoption within the power systems domain.
There are some potential studies for future work. 1) Currently, the dataset we are using is still limited due to accessibility constraints. Thus, building a more comprehensive dataset tailored to a wider range of tasks within the power system domain is possible to improve the performance of FPD; 2) While the current model structure can handle multi-resolution and multi-duration data, it is not yet sufficiently generalized to accommodate data of any resolution or duration. Therefore, there is potential for developing a more generalized structure that can address this challenge and further enhance the versatility of the metric.
Appendix A Appendix: Feature extraction training detail
TABLE V: Regression labels for steady-state training
Regression Tasks
Minimum
Maximum
Range
Variance
Number of data points above average
Number of data points below average
Skewness
Kurtosis
Coefficient of variation
TABLE VI: Hyper-parameters for steady-state model training
Parameter Setting
Minibatch size 1024
Optimizer Adam
Optimizer: learning rate 1e-3
TABLE VII: Hyper-parameters for transient-state model training
Parameter Setting
Minibatch size 128
Optimizer Adam
Optimizer: learning rate 1e-4
Optimizer: weight decay 2e-5
Classification loss coefficient 1.0
Regression loss coefficient 1.0

### v2

@@ IV-E Cross-model comparison with FPD
In the previous sections, all evaluations were conducted directly on the artificially generated datasets. This approach enables precise control over generation quality and model performance assessment. However, FPD can also be used to evaluate model generation performance on the same task, as demonstrated in the following two scenarios.
@@
    Case 1: EV Charging Profile Data Generation
    In [27], the authors present DiffCharge, a diffusion-based generative model that combines an LSTM backbone with multi-head self-attention to conditionally generate EV charging power profiles. They show that DiffCharge better captures short-/long-range temporal dependencies and cross-time interactions than GAN baselines for both battery-level (single-EV sessions) and station-level (aggregate daily load) scenarios. Building on this line of work—and because station-level evaluations were not reported in the underlying study—we train a Wasserstein GAN (W-GAN) baseline on station-level data, compare it against DiffCharge, and assess both using our FPD metric. We also report Kernel distance (RBF) and CRPS. As shown in Fig. 6, DiffCharge better preserves general temporal structure, whereas W-GAN produces unrealistic, long-lasting peaks. Consistently, FPD yields larger distances for W-GAN outputs—aligning with Kernel distance and CRPS—while offering even more apparent discrimination. Hyperparameter choices for DiffCharge and W-GAN appear in SI Sec.S2; Table S5 and Table S7, respectively.
@@
    Case 2: Wind Power Data Generation
    To further demonstrate FPD’s cross-model comparison ability, we replicate the study of [1], which proposed a StyleGAN that integrates conditions on meteorological variables (e.g., temperature, humidity, weather codes) and employs a sequence encoder to infer day-ahead forecast patterns, yielding more realistic generation than GAN baselines such as the conditional W-GAN of [28]. In our replication, we randomly sample day-ahead meteorology paired with the corresponding day-ahead forecast as conditioning inputs and generate 1-day wind-power trajectories (96 15-minute intervals) under a mixed weather scenario. As a baseline, we train a conditional W-GAN with the same conditioning and sampling protocol to generate the same horizon. As shown in Fig. 7, samples from StyleGAN track the overall trend and variability of the real data more closely, whereas W-GAN—despite being meteorology-aware—fails to capture the realistic structure in this case. This ranking is correctly reflected by our proposed FPD metric and by CRPS, while a kernel distance computed directly on the raw data fails to distinguish the models. These results underscore the importance of the feature extraction design in FPD. Hyperparameter choices for StyleGAN and W-GAN appear in SI Sec.S2; Table S6 and Table S7, respectively.
    Figure 6: Case 1 – Comparison of DiffCharge and W-GAN in generating EV charging profiles at the station level.
    Figure 7: Case 2 – Comparison of W-GAN and StyleGAN in generating wind power data.
@@ V Conclusion and Future Work
This paper introduces the Fréchet Power-Scenario Distance (FPD)—a feature-based, distributional metric tailored to smart grid data for evaluating synthetic datasets. FPD encodes both generic temporal structure and domain-specific properties before computing distances in the learned feature space. This design yields consistent, effective judgments across disturbances, time resolutions, and scenario definitions, and better aligns with downstream utility, where raw-data metrics and generic encoders often fall short.
Comprehensive studies demonstrate FPD’s effectiveness in distinguishing data quality across various disturbances, time resolutions, and sub-task scenarios. These results underscore FPD’s potential as a standardized cross-model approach for validating synthetic datasets, offering support for both enhancing model quality and promoting wider adoption within the smart grid domain.
There are several directions for future work. 1) Currently, the dataset we are using is still limited due to accessibility constraints. Thus, building a more comprehensive dataset tailored to a broader range of tasks within the smart grid domain could improve the performance of FPD; 2) While the current model structure can handle multi-resolution and multi-duration data, it is not yet sufficiently general to accommodate data of any resolution or duration. Therefore, there is potential for developing a more generalized structure that can address this challenge and further enhance the versatility of the metric.